b. What is the chance of getting this particular sequence: HHHT? Submit Answer Incorrect. Tries 2/3 Previous Tries c. What is the chance of either getting all heads or no heads on 4 tosses (1/2)A3
Answer #1

here as events are independent and probability of head or tail on a flip =1.2

hence b) probability of HHHT =(1/2)*(1/2)*(1/2)*(1/2) =1/16

c)P(getting all heads or no heds)=P(HHHH)+P(TTTT)=(1/16)+(1/16)=1/8